# Improving End-to-End Training of Retrieval-Augmented Generation Models via Joint Stochastic Approximation

### v1

@@ Abstract
Improving End-to-End Training of Retrieval-Augmented Generation Models via Joint Stochastic Approximation
Hongyu Cao1  Yuxuan Wu1  Yucheng Cai1  Xianyu Zhao2  Zhijian Ou1thanks: Corresponding author. This work is supported by the National Science and Technology Major Project (2023ZD0121401). 1Speech Processing and Machine Intelligence (SPMI) Lab, Tsinghua University, China 2TasiTech Co., Ltd., China ozj@tsinghua.edu.cn
@@ 1 Introduction
Large language models (LLMs) have been shown to store factual knowledge in their parameters through pre-training over large amounts of Internet corpora (Petroni et al., 2019; Brown et al., 2020). However, such implicit knowledge cannot be easily updated, expanded, inspected, and interpreted. Moreover, for many knowledge-intensive tasks, the use of external knowledge beyond the parametric memory of LLMs to generate responses is critical, such as in open-domain question answering (QA) (Chen et al., 2017; Lee et al., 2019; Karpukhin et al., 2020) and knowledge-grouned dialog systems (Kim et al., ; Mishra et al., 2022; Cai et al., 2023). To address these issues, hybrid models that combine parametric memory with nonparametric memories have emerged (Lee et al., 2019; Guu et al., 2020), among which retrieval-augmented generation (RAG) has drawn considerable attention (Lewis et al., 2020).
@@
There have been efforts in developing end-to-end training of an RAG model (Lewis et al., 2020; Zhang et al., 2022; Han et al., 2023; Zamani and Bendersky, 2024), which means eliminating the reliance on intermediate annotations and training all model components simultaneously. In (Lewis et al., 2020), to train the retriever and generator end-to-end, the relevant passage is treated as a discrete latent variable and the following marginal log-likelihood is to be maximized:
@@
Thus end-to-end training of RAG in essence amounts to unsupervised training of a discrete latent-variable model, as shown above. Direct marginalization is intractable; hence, originally, top-K marginalization (TKM) is used for the approximation (Lewis et al., 2020), which we refer to as vanilla RAG. Recently, variational learning (VL) (Kingma and Welling, 2014) has been applied to end-to-end training of RAG in two concurrent and similar works - VRAG (Mishra et al., 2022) and Hindsight (Paranjape et al., ), which we refer to collectively as VRAG. In VRAG, an auxiliary inference model is introduced, acting as a posterior retriever. However, for variational learning of discrete latent variable models, the traditional Monte Carlo gradient estimator for the inference model parameter is known to be either biased or have high-variance (Ou and Song, 2020).
Recently, the joint stochastic approximation (JSA) algorithm (Xu and Ou, 2016; Ou and Song, 2020) has emerged to learn discrete latent variable models with better performance than VL. JSA is a stochastic extension of the EM (expectation-maximization) algorithm and gives unbiased, low-variance stochastic gradients for the inference model.
@@
2 Related work
2.1 Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020) has become a widely recognized paradigm for combining parametric memory with nonparametric memory. A major challenge in end-to-end optimization of RAG models is that the optimization needs to marginalize over relevant passages, which are modeled as discrete latent variables with no annotations. Atlas (Izacard et al., 2023) studies some ad-hoc loss functions (including the vanilla RAG loss via TKM) to train the retriever jointly with generator, and does not observe significant systematic differences between the different training objectives. This highlights the need for more principled end-to-end training method, which our JSA-RAG addresses. In addition to investigating new training methods for RAG, there are other research activities around RAG that are orthogonal to or can benefit from our JSA-RAG. FiD (Izacard and Grave, 2021) presents a new strategy to aggregate and combine multiple passages in decoding. In (Siriwardhana et al., 2023), end-to-end training of RAG is applied to specialized domains such as healthcare and news.
2.2 Learning with discrete latent-variable models
End-to-end training of RAG in essence amounts to learning a discrete latent-variable model. A class of maximum likelihood methods consists of the expectation-maximization (EM) algorithm (Dempster et al., 1977) and its extensions. Variational learning optimizes the Evidence Lower Bound (ELBO) instead of directly maximizing the marginal log-likelihood. VRAG and Hindsight, both based on variational learning, use the TKM approximation to optimize ELBO. RetGen (Zhang et al., 2022) uses the REINFORCE trick (Paisley et al., 2012). Stochastic RAG (Zamani and Bendersky, 2024) uses the Straight-Through trick (Bengio et al., 2013). These parameter estimators are known to be biased or have high-variance (Ou and Song, 2020). The JSA algorithm (Xu and Ou, 2016; Ou and Song, 2020) is a stochastic extension of the EM algorithm with impressive performance, where both the E-step and the M-step (as they cannot be performed exactly) are extended by the stochastic approximation methodology, hence called joint SA. JSA provides us with a new way to improve the end-to-end training of RAG.
@@ 3.2 Training
Training the RAG model from complete data, i.e., knowing $h$, can be easily realized by supervised training. For end-to-end training of the RAG model (i.e., conducting unsupervised training without knowing $h$), we resort to maximizing the marginal likelihood $p_{\theta}(y|x)$ and applying the JSA algorithm (Xu and Ou, 2016; Ou and Song, 2020).
@@
$\mathcal{L}_{\text{JSA}}$ $=-\frac{1}{m}\sum_{i=1}^{m}\left(\log p_{\theta_{r}}(h^{(i)}|x)\right$. (7)
$\left.+\log p_{\theta_{g}}(y|x,h^{(i)})+\log q_{\phi}(h^{(i)}|x,y)\right)$
@@ 3.3 Index rebuilding and passage concatenation
Index Rebuilding.  In previous work, during training, the index of passage embeddings for the KB is often fixed; therefore, the parameters of the passage encoder ($\lambda$) are frozen (Lewis et al., 2020; Mishra et al., 2022; Lin et al., 2023). In this work, to study whether JSA-RAG can perform end-to-end optimization of all modules - including the passage encoder, we explore an index rebuilding scheme. During training, we no longer freeze the parameters of the passage encoder and recalculate the passage embeddings in the index using the updated passage encoder at regular intervals. During passage embedding recalculation, the training process waits; the training is resumed after the index update is completed.
Passage Concatenation.  Note that the prior retriever is improved after end-to-end learning. Inspired by FiD (Izacard and Grave, 2021), we consider a passage concatenation strategy for post-training of the generator while fixing the retriever. The top-$k$ prior retrieved passages are concatenated and append to the context, forming a combined sequence that is fed into the generator for response generation. In this way, the generator is post-trained and in the same way, the generator is used in decoding. This shows the bonus offered by JSA-RAG.
@@ 4.1 Datasets
Open-domain question answering: Unlike ordinary QA, open-domain question answering (ODQA) requires extensive external knowledge to answer questions, which is the primary task explored with RAG systems. We mainly consider three ODQA datasets: NaturalQuestions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), and MS-MARCO (Bajaj et al., 2016)111 Note that the MS-MARCO dataset has two versions, and the version we use in this work is MS-MARCO v1.. For NQ and TQA, we use the Wikipedia December 2018 dump, which contains a total of 24M chunks (passages); for MS-MARCO, instead of using the 10 provided reference passages, we extract its QA pairs and use the MS-MARCO passages from TREC2019 as the KB (Bajaj et al., 2016).
Knowledge grounded dialogs: In dialog datasets, we use conversation history turns as $x$ to retrieve relevant passages and take the response of the current turn as $y$, thus constructing $(x,y)$ pairs. We use the OR-QuAC (Qu et al., 2020) and the DoQA (Campos et al., 2020) datasets. OR-QuAC is an open-domain dialog question answering dataset derived from the QUAC (Question Answering in Context) corpus, requiring models to retrieve and reason over external knowledge to answer multi-turn conversational questions. DoQA comprises of open-ended dialog conversations on different domains like cooking, travel and movies. Both datasets follow the knowledge base settings used in VRAG (Mishra et al., 2022). For OR-QuAC, its KB contains 68k passages, while for DoQA, its KB contains 1.2k passages.
@@ 4.2 Experiment settings
Remarkably, for NQ, TQA, and MS-MARCO without gold passage annotations, the BGE-intialized posterior retriever first retrieves top 100 passages per question from KB, and GPT-4o selects the most relevant as the gold annotation. See Appendix C for details.
@@
Table 1: Generative performance comparison of different models on knowledge-grounded dialog datasets
DoQA OR-QUAC
Method BLEU-4 BLEU-1 F1 BLEU-4 BLEU-1 F1
RAG 15.39 21.69 25.91 6.57 13.51 17.28
VRAG 15.51 21.55 26.02 6.71 13.87 17.63
JSA-RAG 17.11 23.36 27.84 7.76 14.59 18.41
Table 2: Performance comparison of different models on open-domain question answering datasets
NQ TQA MS-MARCO
Method Exact Match Exact Match BLEU-1 Rouge-L
RAG 50.52 72.82 34.23 36.54
VRAG 49.03 72.26 34.14 36.70
JSA-RAG 51.05 75.23 35.28 37.96
@@ 4.3 Main result
Based on the results in Table 1 and Table 2, JSA-RAG demonstrates significant superiority on dialog datasets (DoQA and OR-QUAC) and open-domain QA tasks, with all evaluated metrics outperforming baselines (vanilla RAG and VRAG).
@@
Table 3: Performance evaluation of the prior retrievers for different methods.
Dataset Method R@1 R@10 MRR@10
OR-QUAC DPR 31.16 77.74 48.28
RAG 38.97 83.35 55.93
VRAG 38.91 83.48 55.98
JSA-RAG 39.56 84.76 57.51
DoQA DPR 58.59 83.13 66.94
RAG 67.61 87.33 74.74
VRAG 68.01 87.49 74.78
JSA-RAG 68.09 87.57 75.06
NQ BGE 26.25 63.73 37.64
RAG 27.58 66.97 39.96
VRAG 26.95 64.04 38.21
JSA-RAG 29.23 67.27 41.04
TQA BGE 33.33 66.73 44.21
RAG 36.01 69.48 46.19
VRAG 36.81 70.07 46.73
JSA-RAG 37.39 70.67 46.90
MS-MARCO BGE 10.51 36.81 17.91
RAG 23.17 68.66 36.48
VRAG 23.68 68.81 37.53
JSA-RAG 24.75 71.32 38.65
@@
Figure 2: Comparison of the gradient norms from the posterior retriever for the first 4000 steps during training, using variational and JSA methods respectively.
4.4 Ablations
@@
Variance in gradient norm.  As shown in the Figure 2, we record the gradient norms of the posterior retrievers for JSA-RAG and VRAG every 50 steps during the first 4,000 training steps. The gradient norms of JSA are of low variance, while the gradient norms of VRAG frequently exhibit "sharp" spikes. This observation confirms that during training, JSA-RAG provides gradients with lower variance, enabling more stable training dynamics and thus yielding higher performance.
Table 5: In evaluating the performance of index rebuilding on OR-QuAC, the index is rebuilt every 100 steps during training.
@@
RAG No-rebuild 6.576 38.973 83.353 55.931
Rebuild 8.500 47.772 89.753 65.390
VRAG No-rebuild 6.717 38.910 83.483 55.987
Rebuild 8.583 48.541 90.093 65.901
JSA-RAG No-rebuild 7.760 40.353 84.664 57.970
Rebuild 10.263 49.446 90.115 66.361
4.5 Experiments on index rebuilding and passage concatenation
Index rebuilding.  In the index rebuilding experiment, unlike the main experiments, we do not freeze the parameters of the passage encoder, which is often overlooked in previous work. As Table 5 shows, updating the index every 100 training steps on OR-QuAC improves all metrics for all methods, with substantial gains in retriever performance, and JSA-RAG remains the top performance. These findings demonstrate that incorporating the passage encoder into training and updating the index is effective and that JSA-RAG is able to enhance the performance of all system components in an end-to-end training framework. Meanwhile, it should be noted that rebuilding the knowledge base index for the OR-QuAC dataset takes less than 2 minutes, which hardly affects the training time, since OR-QuAC KB contains 68k passages. We leave asynchronous index rebuilding to work with much larger KBs for future work.
Table 6: In evaluating the passage concatenation method on ODQA, we fix the retrievers and directly concatenate the top 10 retrieved passages with the context, then feed this concatenated input into the generator for post-training and testing.
NQ TQA MS-MARCO
Method concat or not Exact Match Exact Match BLEU-1 Rouge-L
RAG No-concat 50.526 72.828 34.238 36.546
concat 51.105 74.843 34.002 37.319
VRAG No-concat 49.033 72.262 34.148 36.700
concat 51.994 75.541 34.810 37.862
JSA-RAG No-concat 51.053 75.232 35.277 37.961
concat 52.355 76.116 35.613 38.811

### v2

@@ 1 Introduction
Large language models (LLMs) have been shown to store factual knowledge in their parameters through pre-training over large amounts of Internet corpora (Petroni et al., 2019; Brown et al., 2020). However, such implicit knowledge cannot be easily updated, expanded, inspected, and interpreted. Moreover, for many knowledge-intensive tasks, the use of external knowledge beyond the parametric memory of LLMs to generate responses is critical, such as in open-domain question answering (ODQA) (Chen et al., 2017; Lee et al., 2019; Karpukhin et al., 2020) and knowledge-grouned dialog systems (Kim et al., ; Mishra et al., 2022; Cai et al., 2023). To address these issues, hybrid models that combine parametric memory with nonparametric memories have emerged (Lee et al., 2019; Guu et al., 2020), among which retrieval-augmented generation (RAG) has drawn considerable attention (Lewis et al., 2020).
@@
There have been efforts in developing end-to-end training of an RAG model (Lewis et al., 2020; Zhang et al., 2022; Han et al., 2023; Zamani & Bendersky, 2024), which means eliminating the reliance on intermediate annotations and training all model components simultaneously. In (Lewis et al., 2020), to train the retriever and generator end-to-end, the relevant passage is treated as a discrete latent variable and the following marginal log-likelihood is to be maximized:
@@
Thus end-to-end training of RAG in essence amounts to unsupervised training of a discrete latent-variable model, as shown above. Direct marginalization is intractable; hence, originally, top-K marginalization (TKM) is used for the approximation (Lewis et al., 2020), which we refer to as vanilla RAG. Recently, variational learning (VL) (Kingma & Welling, 2014) has been applied to end-to-end training of RAG in two concurrent and similar works - VRAG (Mishra et al., 2022) and Hindsight (Paranjape et al., ), which we refer to collectively as VRAG. In VRAG, an auxiliary inference model is introduced, acting as a posterior retriever. However, for variational learning of discrete latent variable models, the traditional Monte Carlo gradient estimator for the inference model parameter is known to be either biased or have high-variance (Ou & Song, 2020).
Recently, the joint stochastic approximation (JSA) algorithm (Xu & Ou, 2016; Ou & Song, 2020) has emerged to learn discrete latent variable models with better performance than VL. JSA is a stochastic extension of the EM (expectation-maximization) algorithm and gives unbiased, low-variance stochastic gradients for the inference model.
@@ 2.2 Training
Training the RAG model from complete data, i.e., knowing $h$, can be easily realized by supervised training. For end-to-end training of the RAG model (i.e., conducting unsupervised training without knowing $h$), we resort to maximizing the marginal likelihood $p_{\theta}(y|x)$ and applying the JSA algorithm (Xu & Ou, 2016; Ou & Song, 2020).
@@
$\mathcal{L}_{\text{JSA}}$ $=-\frac{1}{m}\sum_{i=1}^{m}\left(\log p_{\theta_{r}}(h^{(i)}|x)\right.\left.+\log p_{\theta_{g}}(y|x,h^{(i)})+\log q_{\phi}(h^{(i)}|x,y)\right)$ (7)
@@ 2.3 Index rebuilding and passage concatenation
Index Rebuilding.  In previous work, during training, the index of passage embeddings for the KB is often fixed; therefore, the parameters of the passage encoder ($\lambda$) are frozen (Lewis et al., 2020; Mishra et al., 2022; Lin et al., 2023a). In this work, to study whether JSA-RAG can perform end-to-end optimization of all modules - including the passage encoder, we explore an index rebuilding scheme. During training, we no longer freeze the parameters of the passage encoder and recalculate the passage embeddings in the index using the updated passage encoder at regular intervals. During passage embedding recalculation, the training process waits; the training is resumed after the index update is completed.
Passage Concatenation.  Note that the prior retriever is improved after end-to-end learning. Inspired by FiD (Izacard & Grave, 2021), we consider a passage concatenation strategy for post-training of the generator while fixing the retriever. The top-$k$ prior retrieved passages are concatenated and append to the context, forming a combined sequence that is fed into the generator for response generation. In this way, the generator is post-trained and in the same way, the generator is used in decoding. This shows the bonus offered by JSA-RAG.
@@ 3.1 Datasets
Open-domain question answering (ODQA) requires extensive external knowledge to answer questions, which is the primary task explored with RAG systems. We mainly consider three ODQA datasets: NaturalQuestions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), and MS-MARCO (Bajaj et al., 2016) [^1] The MS-MARCO dataset has two versions, and the version we use in this work is MS-MARCO v1.. For NQ and TQA, we use the Wikipedia December 2018 dump, which contains a total of 24M chunks (passages); for MS-MARCO, instead of using the 10 provided reference passages, we extract its QA pairs and use the MS-MARCO passages from TREC2019 as the KB (Bajaj et al., 2016).
Knowledge grounded dialogs. In dialog datasets, we use conversation history turns as $x$ to retrieve relevant passages and take the response of the current turn as $y$, thus constructing $(x,y)$ pairs. We use the OR-QuAC (Qu et al., 2020) and the DoQA (Campos et al., 2020) datasets. OR-QuAC is an open-domain dialog question answering dataset derived from the QUAC (Question Answering in Context) corpus, requiring models to retrieve and reason over external knowledge to answer multi-turn conversational questions. DoQA comprises of open-ended dialog conversations on different domains like cooking, travel and movies. Both datasets follow the knowledge base settings used in VRAG (Mishra et al., 2022). For OR-QuAC, its KB contains 68k passages, while for DoQA, its KB contains 1.2k passages.
@@ 3.2 Experiment settings
Remarkably, evaluating retrievers requires access to the gold passages. For NQ, TQA, and MS-MARCO, there are no gold passage annotations. Therefore, we first perform a posterior retrieval based on BGE-large to retrieve the top 100 passages per question from the KB. Then GPT-4o is used to select the most relevant as the gold annotation. So in testing, the retrieval metrics for NQ, TQA, and MS-MARCO are calculated against a high-quality but machine-generated standard. See Appendix C for details. It should also be noted that these GPT-4o-generated annotations are never used in model training, but only used in testing to evaluate the prior retrievers (Table 2) and the posterior retrievers (Table 3), for those datasets without gold passage annotations (i.e., NQ, TQA and MS-MARCO in our experiments).
@@ 3.3 Main result
Based on the results in Table 1, JSA-RAG demonstrates significant superiority on dialog datasets (DoQA and OR-QUAC) and open-domain QA tasks, with all evaluated metrics outperforming baselines (vanilla RAG and VRAG).
@@
Table 1: Performance comparison of different methods on knowledge-grounded dialog and open-domain question answering datasets.
Method Knowledge-Grounded Dialog Open-Domain Question Answering
DoQA OR-QUAC NQ TQA MS-MARCO
BLEU-4 BLEU-1 F1 BLEU-4 BLEU-1 F1 EM EM BLEU-1 Rouge-L
RAG 15.39 21.69 25.91 6.57 13.51 17.28 50.52 72.82 34.23 36.54
VRAG 15.51 21.55 26.02 6.71 13.87 17.63 49.03 72.26 34.14 36.70
JSA-RAG 17.11 23.36 27.84 7.76 14.59 18.41 51.05 75.23 35.28 37.96
Table 2: Performance evaluation of the prior retrievers for different methods. The base retrievers for knowledge-grounded dialog datasets and open-domain question answering datasets are BGE-large and DPR, respectively.
Method Knowledge-Grounded Dialog Open-Domain Question Answering
OR-QUAC DoQA NQ TQA MS-MARCO
R@1 R@10 MRR@10 R@1 R@10 MRR@10 R@1 R@10 MRR@10 R@1 R@10 MRR@10 R@1 R@10 MRR@10
Base 31.16 77.74 48.28 58.59 83.13 66.94 26.25 63.73 37.64 33.33 66.73 44.21 10.51 36.81 17.91
RAG 38.97 83.35 55.93 67.61 87.33 74.74 27.58 66.97 39.96 36.01 69.48 46.19 23.17 68.66 36.48
VRAG 38.91 83.48 55.98 68.01 87.49 74.78 26.95 64.04 38.21 36.81 70.07 46.73 23.68 68.81 37.53
JSA-RAG 39.56 84.76 57.51 68.09 87.57 75.06 29.23 67.27 41.04 37.39 70.67 46.90 24.75 71.32 38.65
3.4 Ablations
@@
Variance in gradient norm. As shown in Figure 2, we record the gradient norms of the posterior retrievers for JSA-RAG and VRAG every 50 steps during the first 4,000 training steps. The gradient norms of JSA are of low variance, while the gradient norms of VRAG frequently exhibit ”sharp” spikes. This observation confirms that during training, JSA-RAG provides gradients with lower variance, enabling more stable training dynamics and thus yielding higher performance.
Figure 2: Comparison of the gradient norms from the posterior retriever.
3.5 Experiments on index rebuilding and passage concatenation
Index rebuilding.  In the index rebuilding experiment, unlike the main experiment, we do not freeze the parameters of the passage encoder. Every 100 steps, we recalculate the passage embeddings using the updated passage encoder and rebuild the index. As Table 5 shows, updating the index on OR-QuAC improves all metrics for all methods, with substantial gains in retriever performance, and JSA-RAG remains the top performance. These findings show that JSA-RAG is able to enhance the performance of all system components in an end-to-end training framework. Meanwhile, it should be noted that in our experiments, model training and index rebuilding are serially conducted. In the OR-QuAC (68k passages) experiment, it takes about 3 minutes for 100 training steps, and then index rebuilding takes less than 2 minutes, which hardly affects the training time. However, for larger scale experiments with Wikipedia (24M packages), index rebuilding takes 6 hours with the hardware in our experiments (4*A100 (40G)). When we used a long interval for index rebuiding (we tried 2000-5000 steps in experiments), marginal performance improvements were obtained compared to the results of not using index rebuilding, since the model training behavior does not change much. However, if we reduce the interval for index rebuilding (e.g., every 100 steps), then it will add 200×6 hours with huge time cost (for a total of 20,000 steps for one experiment). Asynchronous algorithm research (parameter training and index rebuilding are performed in parallel) is interesting future work. It should be noted that such asynchronous algorithm is orthogonal to the core contribution of this work – the developement of JSA-RAG beyond the prior arts which are vanilla RAG (using top-K marginalization) and variational approaches like VRAG.
Table 4: Performance of index rebuilding on OR-QuAC (index rebuilt every 100 steps during training).
@@ 3.5 Experiments on index rebuilding and passage concatenation
RAG 6.57 38.97 83.35 55.93
8.50 47.77 89.75 65.39
VRAG 6.71 38.91 83.48 55.98
8.58 48.54 90.09 65.90
JSA-RAG 7.76 40.35 84.66 57.97
10.26 49.44 90.11 66.36
Table 5: Performance of passage concatenation on ODQA (top 10 passages concatenated with context).
Method Concat NQ TQA MS-MARCO
EM EM BLEU-1 Rouge-L
RAG 50.52 72.82 34.23 36.54
51.10 74.84 34.00 37.31
VRAG 49.03 72.26 34.14 36.70
51.99 75.54 34.81 37.86
JSA-RAG 51.05 75.23 35.27 37.96
52.35 76.11 35.61 38.81
4 Related work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) has become a widely recognized paradigm for combining parametric memory with nonparametric memory. A major challenge in end-to-end optimization of RAG models is that the optimization needs to marginalize over relevant passages, which are modeled as discrete latent variables with no annotations. Atlas (Izacard et al., 2023) studies some ad-hoc loss functions (including the vanilla RAG loss via TKM) to train the retriever jointly with generator, and does not observe significant systematic differences between the different training objectives. This highlights the need for more principled end-to-end training method, which our JSA-RAG addresses. In addition to investigating new training methods for RAG, there are other research activities around RAG. FiD (Izacard & Grave, 2021) presents a new strategy to aggregate and combine multiple passages in decoding. In (Siriwardhana et al., 2023), end-to-end training of RAG is applied to specialized domains such as healthcare and news. CoV-RAG (He et al., 2024) integrates a verification module into RAG and uses verification data to finetune RAG generators, while keeping retrievers frozen. RAT (Wang et al., 2024) combines RAG with chain of thought (CoT) prompting but does not involve any model training. RA-DIT (Lin et al., 2023b) finetunes retrievers and generators separately on a set of multi-task instruction-tuning datasets, which are not end-to-end optimized. These recent works are orthogonal to and can benefit from JSA-RAG, which focuses on improving end-to-end training of RAG models.
Learning with discrete latent-variable models.  End-to-end training of RAG in essence amounts to learning a discrete latent-variable model. A class of maximum likelihood methods consists of the expectation-maximization (EM) algorithm (Dempster et al., 1977) and its extensions. Variational learning optimizes the Evidence Lower Bound (ELBO) instead of directly maximizing the marginal log-likelihood. VRAG and Hindsight, both based on variational learning, use the TKM approximation to optimize ELBO. RetGen (Zhang et al., 2022) uses the REINFORCE trick (Paisley et al., 2012). Stochastic RAG (Zamani & Bendersky, 2024) uses the Straight-Through trick (Bengio et al., 2013). These parameter estimators are known to be biased or have high-variance (Ou & Song, 2020). The JSA algorithm (Xu & Ou, 2016; Ou & Song, 2020) is a stochastic extension of the EM algorithm with impressive performance, where both the E-step and the M-step (as they cannot be performed exactly) are extended by the stochastic approximation methodology, hence called joint SA. To the best of our knowledge, our work is the first to apply and implement the general JSA to successfully realize more powerful and principled end-to-end training method for RAG.